# On De Giorgi’s Conjecture of Nonlocal approximations for free-discontinuity problems: The symmetric gradient case

### v2

@@ Abstract.
Keywords and phrases: Free-discontinuity problems, nonlocal approximations, symmetric gradient, integral geometric measures.
@@ 1. Introduction
An important remark concerns the fact that the energies in (1.1) only approximate the Mumford-Shah functional under suitable additional assumptions on the geometry of $\Omega$, such as Lipschitz regularity (see [31, Remark 7.1]). Such requirements may not be fully consistent when handling free-boundary problems where discontinuities are already present within the material. Consider, for example, the case of a crack-initiated domain, where the presence of an $(n-1)$-dimensional set $\Gamma\subset\Omega$, representing an initial crack makes the set $\Omega\setminus\Gamma$ not Lipschitz.
@@
The fact that (1.8) is replaced in our setting by (1.9) is a huge obstacle to establish the structural properties of the limiting function space. In particular, the technique presented in [26] for deriving the GBD-structural properties fails, and we need to rely on the more recent integral geometric approach introduced in [2]. The main difference from the approach in [2] lies in the role of the flatness of $\mathbb{R}^{n}$, where rescaling and translating the domain preserve straight lines as geodesics, maintaining invariance in the associated GBD-space. Furthermore, the ${\rm GDB}$-theory developed in [25] provides a quite general method to relate the jump set of $u$ to the behavior of codimension one slices, which we are able to adapt to our setting. We remark that such an approach seems not to be successful in a Riemannian manifold. In Propositions 3.143.16, we leverage the above features to avoid relying on a Korn-Poincaré-type inequality to estimate the dimensionality of the set on which the jump of all one-dimensional slices of $u$ concentrate. Ultimately, we conclude in Theorem 3.1 that the limit map $u$ belongs to $\text{GSBD}(\Omega)$, and thus characterize the domain of the $\Gamma$-limit of $\mathcal{F}^{p}_{\varepsilon}$. For the detailed definitions of the relevant quantities and the proof of the structural properties, we refer the reader to Sections 2 and 3, respectively.
An alternative approach. After this paper was submitted, the authors became aware that another equivalent characterization of $GSBD$, based on controlling the measures $\hat{\mu}^{\xi}_{u}$ through finitely many direction $\xi$, has been established in [21]. Instead we adopt here a totally different perspective, by first deriving fine properties of the limiting space and then proving it coincides with $GSBD$ via some integro-geometrical arguments.
@@ 2. Definitions and notation
$\hat{\mu}^{p}_{u}(U):=\sup_{\mathscr{B}}\sum_{B\in\mathscr{B}}\bigg{(}\int_{\mathbb{S}^{n-1}}\hat{\mu}^{\xi}_{u}(B)^{p}\,\mathrm{d}\mathcal{H}^{n-1}(\xi)\bigg{)}^{\frac{1}{p}}\qquad\text{for $U\subset\Omega$ open,}$, (2.3)
@@
[…]
@@ 3.1. Preliminaries
We start with a technical proposition for one-dimensional functions. We introduce the class $\mathcal{T}$ of all functions $\tau\in C^{1}(\mathbb{R})$ such that $-\frac{1}{2}\leq\tau\leq\frac{1}{2}$ and $0\leq\tau^{\prime}\leq 1$. We recall that, given an open set $I\subset\mathbb{R}$ and given $(\tau_{j})\subset C^{1}(I)$, $\tau\in C^{1}(I)$, we have $\tau_{j}\to\tau$ in $C^{1}_{loc}(I)$ if and only if for every open set $U\Subset I$ it holds $\|\tau_{j}-\tau\|_{L^{\infty}(U)}+\|\tau^{\prime}_{j}-\tau^{\prime}\|_{L^{\infty}(U)}\to 0$ as $j\to\infty$.
@@
By combining Theorem 3.6 together with the countably $(n-1)$-rectifiability of $J_{u}$ (see [27]), we make use of the Area Formula to infer for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ that for every Borel set $B\subset\Omega$
@@
where $D^{j}$ denotes the jump part of the distributional derivative. In particular, there exists a dimensional constant $10<c_{n}<$ 1 such that for every $v\in\mathbb{R}^{n}$ and $\nu\in\mathbb{S}^{n-1}$
@@
[…]
@@
Thanks to Proposition 3.11, we are now in a position to identify a precise representative for co-dimension one slices through $(n-1)$-planes.
@@
[…]
@@
Assume that Theorem 3.6 holds true up to dimension $n-1$. Let $u\colon\Omega\to\mathbb{R}^{n}$ be measurable. Then, it holds
@@
Indeed, from the definition of $\text{GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$, we know that for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ we have $\hat{u}^{\xi}_{y}\in\text{BV}_{loc}(\Omega^{\xi}_{y})$ for $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$. Equivalently, for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ we have $\hat{u}^{\xi}_{\pi_{\xi}(x)}\in\text{BV}_{loc}(\Omega^{\xi}_{\pi_{\xi}(x)})$ for $\mathcal{L}^{n}$-a.e. $x\in\Omega$, where $\pi_{\xi}\colon\mathbb{R}^{n}\to\Pi^{\xi}$ denotes the orthogonal projection. Therefore, by using Fubini’s Theorem on the product space $\Omega\times\mathbb{S}^{n-1}$ with measure $\mathcal{L}^{n}\otimes\mathcal{H}^{n-1}$, and the fact that $\hat{u}^{\xi}_{x}=\hat{u}^{\xi}_{\pi_{\xi}(x)}$, we immediately infer the validity of (3.21). To prove (3.23), in view of the measurability (3.22) (see Lemma 2.1), it is enough to make use of Fubini’s Theorem and exchange the order of integration in $x$ and $\xi$.
@@
Let us assume by contradiction that the measure $\hat{\mu}^{1}_{u}$ has null $(n-1)$-density at $x$ while the set $\Sigma\subset\mathbb{S}^{n-1}$ defined as
@@
We proceed by induction. Observe that in dimension one the theorem is trivially true. Now, assume that the theorem holds true in dimension $n-1$. Therefore, from Proposition 3.14 we infer that the measure $\mathscr{I}_{u,1}$ concentrates on the complement in $\Omega$ of the set $E$ defined in (3.18). Since by Proposition 3.16 we have that $\Omega\setminus E\subset\{x\in\Omega:\Theta^{*n-1}(\hat{\mu}^{1}_{u},x)>0\}$, then by applying the classical density estimates for Radon measure (cf. [29]) we infer that $\Omega\setminus E$ is $\sigma$-finite with respect to $\mathcal{H}^{n-1}$. We are thus in position to apply Besicovitch-Federer structure theorem [29] to write $\Omega\setminus E=R\cup U$ where $R\subset\Omega$ is countably $(n-1)$-rectifiable while the Borel set $U\subset\Omega$ satisfies
@@ 4.1. Preliminary results
$\tau(u_{k}\cdot\eta_{i})\to f^{i}_{\infty}\qquad\text{pointwise a.e.~{}in $\Omega$ for every $i=1,\dotsc,n$ as $k\to\infty$},$
@@
$(\hat{u}_{k})_{y,\lambda}^{\xi}\to\hat{u}_{y,\lambda}^{\xi}\qquad\text{in $L^{1}_{loc}((\Omega\setminus A)_{y}^{\xi})$ for every $\xi$ and for $\mathcal{H}^{n}$-a.e.~{}$y\in\Pi^{\xi}$},$ (4.41)
@@ 5.1. $\Gamma$-convergence
Since $I$ is an interval, the set $I\cap(I-\varepsilon)$ is also an interval or it is empty. Assume that it is not empty otherwise there is nothing to prove. Define $A^{\varepsilon}:=\{t\in I\cap(I-\varepsilon):[t,t+\varepsilon]\cap J_{u}\neq\emptyset\}$. Notice that $t\in I\cap(I-\varepsilon)$ implies $[t,t+\varepsilon]\subset I$. We have
@@ 5.2. Convergence of quasi-minimisers
and that the term $\mathcal{H}^{n-1}(J_{u}\cap\partial_{D}\Omega)$ penalizes the part of $\partial_{D}\Omega$ on which the Dirichlet condition is not attained, namely, the set $\{x\in\partial_{D}\Omega:u^{-}(x)\neq f(x)\}$ where $u^{-}(x)$ denotes the trace with respect to the inner normal to $\partial\Omega$ of $u$ at $x$

### v3

@@ Abstract.
Keywords and phrases:GSBD Free-discontinuity problems, nonlocal approximations, symmetric gradient, integral geometric measures.
@@ 1. Introduction
An important remark concerns the fact that the energies in (1.1) only approximate the Mumford-Shah functional under suitable additional assumptions on the geometry of $\Omega$, such as Lipschitz regularity (see [31, Remark 7.1]). Such requirements may not be fully consistent when handling free-boundary problems where discontinuities are already present within the material. Consider, for example, the case of a crack-initiated domain, where the presence of an (n1)(n-1)-dimensional set $\Gamma\subset\Omega$, representing an initial crack makes the set $\Omega\setminus\Gamma$ not Lipschitz.
@@
The fact that (1.8) is replaced in our setting by (1.9) is a huge obstacle to establish the structural properties of the limiting function space. In particular, the technique presented in [26] for deriving the GBD-structural properties fails, and we need to rely on the more recent integral geometric approach introduced in [2]. The main difference from the approach in [2] lies in the role of the flatness of $\mathbb{R}^{n}$, where rescaling and translating the domain preserve straight lines as geodesics, maintaining invariance in the associated GBD-space. Furthermore, the $\rm GDB$}-theory developed in [25] provides a quite general method to relate the jump set of $u$ to the behavior of codimension one slices, which we are able to adapt to our setting. We remark that such an approach seems not to be successful in a Riemannian manifold. In Propositions 3.143.16, we leverage the above features to avoid relying on a Korn-Poincaré-type inequality to estimate the dimensionality of the set on which the jump of all one-dimensional slices of $u$ concentrate. Ultimately, we conclude in Theorem 3.1 that the limit map $u$ belongs to $\text{GSBD}(\Omega)$, and thus characterize the domain of the $\Gamma$-limit of $\mathcal{F}^{p}_{\varepsilon}$. It is worth noting that, although it is not used in the present paper, this method also allows one to address the case of jump sets with infinite length, which may be of interest for cohesive models obtained as the $\Gamma$-limit of finite-difference approximating functionals, as in [32]. For the detailed definitions of the relevant quantities and the proof of the structural properties, we refer the reader to Sections 2 and 3, respectively.
An alternative approach.
After the first version of this paper was submitted as a Preprint on arXiv, a new characterization of GSBD with finite jump was proved in [21], from which it was deduced that our limiting space is GSBD. This characterization was based on controlling the measures $\hat{\mu}^{\xi}_{u}$ through finitely many direction $\xi$, which turns out to be equivalent to a control on average. In this new version we propose a totally different perspective, by first deriving fine properties of the limiting space and then proving it coincides with GSBD via some integro-geometrical arguments, hence providing an alternative proof to [21].
@@ 2. Definitions and notation
$\hat{\mu}^{p}_{u}(U):=\sup_{\mathscr{B}}\sum_{B\in\mathscr{B}}\bigg(\int_{\mathbb{S}^{n-1}}\hat{\mu}^{\xi}_{u}(B)^{p}\,\mathrm{d}\mathcal{H}^{n-1}(\xi)\bigg)^{\frac{1}{p}}\qquad\text{for $U\subset\Omega$ open,}$ (2.3)
@@
$\mathcal{F}^{\prime p}_{\varepsilon}(u,\Omega):=\sup_{\mathscr{B}}\sum_{B\in\mathscr{B}}\bigg(\int_{\frac{B-B}{\varepsilon}}F_{\varepsilon,\xi}(u,B)^{p}e^{-|\xi|^{2}}\,d\xi\bigg)^{\frac{1}{p}}$
@@ 3.1. Preliminaries
We start with a technical proposition for one-dimensional functions. We introduce the class $\mathcal{T}$ of all functions $\tau\in C^{1}(\mathbb{R})$ such that $\frac{1}{2}\leq\tau\leq\frac{1}{2}$ and $0\leq\tau^{\prime}\leq 1$. We recall that, given an open set $I\subset\mathbb{R}$ and given $(\tau_{j})\subset C^{1}(I)$, $\tau\in C^{1}(I)$, we have $\tau_{j}\to\tau$ in $C^{1}_{loc}(I)$ if and only if for every open set $U\Subset I$ it holds $\|\tau_{j}-\tau\|_{L^{\infty}(U)}+\|\tau^{\prime}_{j}-\tau^{\prime}\|_{L^{\infty}(U)}\to 0$ as $j\to\infty$.
@@
By combining Theorem 3.6 together with the countably (n1)(n-1)-rectifiability of $J_{u}$ (see [27]), we make use of the Area Formula to infer for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ that for every Borel set $B\subset\Omega$
@@
where $D^{j}$ denotes the jump part of the distributional derivative. In particular, there exists a dimensional constant $0<c_{n}<1$ such that for every $v\in\mathbb{R}^{n}$ and $\nu\in\mathbb{S}^{n-1}$
@@
$y\mapsto\sum_{t\in B^{\xi}_{y}}(|[\hat{u}^{\xi}_{y}(t)]|\wedge 1)\qquad\text{ is $\mathcal{H}^{n-1}$-measurable}$ (3.11)
$\xi\mapsto\int_{\Pi^{\xi}}\sum_{t\in B^{\xi}_{y}}\big(|[\hat{u}^{\xi}_{y}(t)]|\wedge 1\big)\,\mathrm{d}\mathcal{H}^{n-1}(y)\qquad\text{ is $\mathcal{H}^{n-1}$-measurable}.$ (3.12)
@@
Thanks to Proposition 3.11, we are now in a position to identify a precise representative for co-dimension one slices through (n1)(n-1)-planes.
@@
First of all we notice that the Borel regular measure on $\mathbb{S}^{n-1}$ defined as $B\mapsto\int_{\mathbb{S}^{n-1}}\mathcal{H}^{n-2}(B\cap\Pi^{\xi})\,d\mathcal{H}^{n-1}(\xi)$ is invariant under the action of the orthogonal group. Hence from [29, Theorem 2.7.7] we infer the existence of a dimensional constant $c(n)$ such that
$\mathcal{H}^{n-1}(B)=c(n)\int_{\mathbb{S}^{n-1}}\mathcal{H}^{n-2}(B\cap\Pi^{\xi})\,d\mathcal{H}^{n-1}(\xi)\quad\forall B\subset\mathbb{S}^{n-1}\text{ Borel set},$ (3.16)
@@
Assume that Theorem 3.6 holds true up to dimension n1n-1. Let $u\colon\Omega\to\mathbb{R}^{n}$ be measurable. Then, it holds
@@
$\ \ \ \ \ \ (x,\xi)\mapsto O^{\xi}_{x}(u)\ \ \text{ is $(\mathcal{L}^{n}\otimes\mathcal{H}^{n-1})$-measurable}$ (3.22)
@@
Let us assume by contradiction that the measure $\hat{\mu}^{1}_{u}$ has null (n1)(n-1)-density at $x$ while the set $\Sigma\subset\mathbb{S}^{n-1}$ defined as
@@
We proceed by induction. Observe that in dimension one the theorem is trivially true. Now, assume that the theorem holds true in dimension n1n-1. Therefore, from Proposition 3.14 we infer that the measure $\mathscr{I}_{u,1}$ concentrates on the complement in $\Omega$ of the set $E$ defined in (3.18). Since by Proposition 3.16 we have that $\Omega\setminus E\subset\{x\in\Omega:\Theta^{*n-1}(\hat{\mu}^{1}_{u},x)>0\}$, then by applying the classical density estimates for Radon measure (cf. [29]) we infer that $\Omega\setminus E$ is $\sigma$-finite with respect to $\mathcal{H}^{n-1}$. We are thus in position to apply Besicovitch-Federer structure theorem [29] to write $\Omega\setminus E=R\cup U$ where $R\subset\Omega$ is countably (n1)(n-1)-rectifiable while the Borel set $U\subset\Omega$ satisfies
@@ 4.1. Preliminary results
[…]
@@
$|(\hat{u}_{k})_{y,\lambda}^{\xi}|\to|\hat{u}_{y,\lambda}^{\xi}|=\lambda\qquad\text{in $L^{1}_{loc}(A_{y}^{\xi})$ for a.e.~$\xi$, for $\mathcal{H}^{n-1}$-a.e.~$y\in\Pi^{\xi}$,}$ (4.42)
@@ 5.1. $\Gamma$-convergence
Since $I$ is an interval, the set $I\cap(I-\varepsilon)$ is also an interval or it is empty. Assume that it is not empty otherwise there is nothing to prove. Define $A^{\varepsilon}:=\{t\in I\cap(I-\varepsilon):[t,t+\varepsilon]\cap J_{u}\neq\emptyset\}$. Notice that $t\in I\cap(I-\varepsilon)$ implies $t,t+\varepsilon]\subset I$. We have
@@ 5.2. Convergence of quasi-minimisers
and that the term $\mathcal{H}^{n-1}(J_{u}\cap\partial_{D}\Omega)$ penalizes the part of $\partial_{D}\Omega$ on which the Dirichlet condition is not attained, namely, the set $\{x\in\partial_{D}\Omega:u^{-}(x)\neq f(x)\}$ where $u^{-}(x)$ denotes the trace with respect to the inner normal to $\partial\Omega$ of $u$ at $x$ .
@@ Acknowledgements
We would like to thank Antonin Chambolle and Vito Crismale for pointing out to us, after the first version of this paper was written, that the identification of the spaces $GSBD$ and $GSBV^{\mathcal{E}}$ was indeed possible. This observation encouraged us to develop an alternative approach.